OpenStudy (logan13):

A 12-foot ladder is leaning against a wall. The bottom of the ladder is 5 feet away from the bottom of the wall. Approximately how high up the wall does the top of the ladder reach? A. 2.4 feet B. 7.0 feet C. 13.0 feet D. 10.9 feet

OpenStudy (welshfella):



OpenStudy (welshfella):

use the Pythagoras theorem 12^2 = 5^2 + x^2 solve for x

OpenStudy (logan13):

so B

OpenStudy (logan13):

@paki

OpenStudy (welshfella):

x^2 = 12^2 - 5^2 can you work that out use your calculator

OpenStudy (mrcoolguy):

12-5 = 7 so B